Suppose the coefficient of kinetic friction betwee
Answer #1

a) T −mAgsinθ−μk*FN =mAaA ............. (1)

T=mBg−mBa ...........(2)

From (1) and (2),
mBg−mBa−mAgsinθ−μk*mAgcosθ=mAaA

g−a−gsinθ−μk*gcosθ= a

10 - a - 5.3 - 1.018 = a
a = 1.84 m/s2


b) To keep it stationary,
mg = mg(sinθ + μk*cosθ)
0.53 + μk*0.848 = 1
μk = 0.554